OpenStudy (anonymous):

x+2 + (x-1)(x+2) --------- ---------- (x-3)(x+1) x-3 simplify the sum.

OpenStudy (campbell_st):

ok... so you are being asked to add 2 fractions.... so the same rules apply for numeric factions apply. 1. you need a common denominator which appears to be (x -3)(x + 1) so the 2nd fraction needs (x+1) in the numerator and denominator so you have \[\frac{(x +2)}{(x -3)(x +1)} + \frac{(x -1)(x +2)(x +1)}{(x -3)(x + 1)}\] so you can write it as \[\frac{(x +2)(1 + (x-1)(x+1))}{(x -3)(x +1)}\] I'll let you simplify ti further.

OpenStudy (anonymous):

I ended up getting... \[\frac{ x ^{3}+2x ^{2} }{ (x-3)(x+1) }\] i told you i am bad at this...

OpenStudy (campbell_st):

yep. that looks good, you may need to just distribute the denominator... I'm not sure what form you answer needs to be in.

OpenStudy (anonymous):

you leave the denominator as is! thank you so much, HUGE HELP!